P is of degree 3; P(0)=-24; zeros: 3,4,2
what is the polynomial function in standard form?

Answer by lwsshak3(11628)

P is of degree 3; P(0)=-24; zeros: 3,4,2
what is the polynomial function in standard form?
..
(x-3)(x-4)(x-2)
=(x^2-7x+12)(x-2)
=x^3-7x^2+12x-2x^2+14x-24
=x^3-9x^2+26x-24
P(x)=x^3-9x^2+26x-24